# Jinhang Chai     Jianqing Fan

### v1

@@ Appendix B Key Separation Lemma
B.1 Proof sketch
We first prove a simple discrepancy lemma B.2. Then we do more refined discrepancy analysis in Lemma B.3 and B.4 in the sense of norm discrepancy and direction discrepancy respectively. Combine them we get Lemma B.5 which is a special case of Lemma B.1 in which every singular values are the same. We then tackle the general case utilizing a technical lemma B.6 which is proved using linear programming technique.
Lemma B.2.
@@ B.1 Proof sketch
Lemma B.3.
Let $p>r\geq 2$, $\ U,V,L,R\in\mathcal{O}_{p,r}$, and assume they are all $\mu$-incoherent, denote $u=U^{\top}e_{1}$ and similarly for $v,l,r$. Suppose $u=ae_{1}$ and $l=be_{1}$, $a\leq b$ then we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{8\mu r}\cdot r_{1}^{2}(a-b)^{2}$
Lemma B.4.
Under the same assumptions of B.3, we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{32\mu r^{3}}a^{2}(r_{1}-v_{1})^{2}$
Combining the previous two lemmas B.3 and B.4 gives the next lemma which characterizes how the discrepancy between the first rows of $U,V,L,R$ impact the inner product $\langle U^{\top}L,V^{\top}R\rangle$.
Lemma B.5.
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq c\frac{p}{\mu r^{3}}\cdot(u^{\top}v-l^{\top}r)^{2}$
Lemma B.5 solves the case where all singular values are identical for both $P$ and $Q$. Now we utilize the special case to prove the general case. We pause for a bit and state a technical lemma utilizing linear programming.
Lemma B.6.
Consider the optimization problem which is essentially an LP problem.
@@
$\Omega=\Big{\{}X\in\mathbb{R}^{r\times r}\Big{|}\sum_{i=1}^{r}|X_{i,k}|\leq 1,\ \sum_{i=1}^{r}|X_{k,i}|\leq 1,\ \sum_{i=1}^{k}\sum_{j=1}^{k}X_{i,j}\leq k-\gamma_{k}\quad\forall 1\leq k\leq r\Big{\}}$ (31)
@@
$x_{i}=(I-e_{1}e_{1}^{\top})Ue_{i},\quad y_{i}=(I-e_{1}e_{1}^{\top})Le_{i}$
@@
$U=\begin{bmatrix}a&0&\cdots&0\\ x_{1}&x_{2}&\cdots&x_{r}\end{bmatrix},\quad L=\begin{bmatrix}b&0&\cdots&0\\ y_{1}&y_{2}&\cdots&y_{r}\end{bmatrix},$
then
@@
$\begin{bmatrix}\frac{x_{1}}{\sqrt{1-a^{2}}},&x_{2},&x_{3},&\cdots,&x_{r}\end{bmatrix}\in\mathcal{O}_{p-1,r}$
In other words, they form a incomplete basis of $\mathbb{R}^{p-1}$, hence
$(\frac{x_{1}^{\top}y_{1}}{\sqrt{1-a^{2}}})^{2}+\sum_{i=2}^{r}(x_{i}^{\top}y_{1})^{2}\leq\|y_{1}\|^{2}=1-b^{2}$ (34)
@@
$\mathcal{E}_{r}=\Bigg{\{}\begin{bmatrix}s_{1}\\ s_{2}\\ \cdots\\ s_{r}\end{bmatrix}\in\mathbb{R}^{r}\Bigg{|}\frac{(s_{1}-ab)^{2}}{1-a^{2}}+\sum_{i=2}^{r}s_{i}^{2}\leq 1-b^{2}\Bigg{\}}$
@@
Now we turn to $V^{\top}R$, it turns out it suffices to bound $e_{1}^{\top}V^{\top}Re_{1}$ to prove this lemma. In fact, using Lemma B.2 with $F=Ve_{1},G=Re_{1},r=1$, we get
$|e_{1}^{\top}V^{\top}Re_{1}|=\|F^{\top}G\|_{F}\leq\sqrt{1-(v_{1}-r_{1})^{2}}\leq 1-\frac{(v_{1}-r_{1})^{2}}{2}$ (35)
@@
It turns out the structures of $\mathcal{E}_{r}$ and $\mathcal{T}_{r}$ is enough to prove the lemma. Specifically, we can prove for all $s\in\mathcal{E}_{r}$ and $t\in\mathcal{T}_{r}$, it holds that
$1-s^{\top}t\geq\frac{p}{8\mu r}\cdot r_{1}^{2}(a-b)^{2}$ (36)
@@
$\begin{bmatrix}\widetilde{s}_{1}\\ \widetilde{s}_{2}\end{bmatrix}\in\mathcal{E}_{2},\begin{bmatrix}\tilde{t}_{1}\\ \tilde{t}_{2}\end{bmatrix}\in\mathcal{T}_{2}$ (38)
while by Cauchy’s Inequality, $s^{\top}t\leq\widetilde{s}_{1}\tilde{t}_{1}+\widetilde{s}_{2}\tilde{t}_{2}=\widetilde{s}^{\top}\tilde{t}$. Therefore we reduce to the case $r=2$.
@@
In fact, it suffices to rule out the case that the optimal $t^{\star}$ lies on the vertical line part of the border of $\mathcal{T}_{2}$. Suppose without loss of generality that $|t^{\star}_{1}|=1-\frac{(v_{1}-r_{1})^{2}}{2},\ t^{\star}_{2}\geq 0$, we immediate have the optimal $s^{\star}_{2}\geq 0$. Now define $\tilde{t}:=(t^{\star}_{1},\sqrt{1-t_{1}^{*2}})$. Let $\delta_{t}=\sqrt{1-t_{1}^{*2}}-t^{\star}_{2}>0$. We have $\tilde{t}\in\mathcal{T}_{2}$ and $s^{*T}\tilde{t}=s^{*T}t^{\star}+s^{\star}_{2}\delta_{t}>s^{*T}t^{\star}$, which is in contradiction with the optimality of $(s^{\star},t^{\star})$. Therefore, $t^{\star}$ do take the form of $\begin{bmatrix}\cos(\phi)\\ \sin(\phi)\end{bmatrix}$.
Let $\lambda:=|\tan(\phi)|$. As $t^{\star}$ lies on the border of $\mathcal{T}$ other than the vertical line $|t_{1}|=1-\frac{(v_{1}-r_{1})^{2}}{2}$, we have $\cos(\phi)\leq 1-\frac{(v_{1}-r_{1})^{2}}{2}$. Notice by the $\mu$-incoherence condition, $|v_{1}|\leq\sqrt{\frac{\mu r}{p}},|r_{1}|\leq\sqrt{\frac{\mu r}{p}}$,
$(v_{1}-r_{1})^{2}\leq 4(v_{1}^{2}+r_{1}^{2})\leq 4\frac{\mu r}{p}$
An observation here is that we can assume $\frac{\mu r}{p}\leq\frac{1}{8}$ when we take $\tilde{c}\geq 8$ in Lemma B.1. Otherwise, $\frac{\tilde{c}\mu r^{4}}{p}\geq\frac{8\mu r}{p}\geq 1$ and Lemma B.1 is vacuous. Assuming $\frac{\mu r}{p}\leq\frac{1}{8}$, it holds that $\cos(\phi)\leq 1-\frac{1}{4}\leq\frac{1}{\sqrt{2}}$. Therefore, $\lambda=|\tan(\phi)|\leq 1$.
As $s^{\star}$ lies on the ellipse, it takes the form of $\begin{bmatrix}ab+\sqrt{(1-a^{2})(1-b^{2})}\cos{\theta}\\ \sqrt{1-b^{2}}\sin{\theta}\end{bmatrix}$.
Therefore we have
@@
$=\frac{ab+\sqrt{(1-a^{2})(1-b^{2})}\cos{\theta}+\lambda\sqrt{1-b^{2}}\sin{\theta}}{\sqrt{1+\lambda^{2}}}$
$\leq\frac{ab+\sqrt{(1-a^{2})(1-b^{2})+\lambda^{2}(1-b^{2})}}{\sqrt{1+\lambda^{2}}}$
Observe that
$1-\frac{ab+\sqrt{(1-a^{2})(1-b^{2})+\lambda^{2}(1-b^{2})}}{\sqrt{1+\lambda^{2}}}=E_{1}+E_{2}$
where
$E_{1}$ $=\frac{\big{(}a^{2}+\sqrt{(1-a^{2})^{2}+\lambda^{2}(1-a^{2})}\big{)}-\big{(}ab+\sqrt{(1-a^{2})(1-b^{2})+\lambda^{2}(1-b^{2})}\big{)}}{\sqrt{1+\lambda^{2}}}$
$=\frac{\big{(}a^{2}-ab\big{)}+\sqrt{(1-a^{2}+\lambda^{2})}\big{(}\sqrt{1-a^{2}}-\sqrt{1-b^{2}}\big{)}}{\sqrt{1+\lambda^{2}}}$
Now notice that the numerator is increasing with $\lambda^{2}$ and $\lambda^{2}\leq 1$, it holds that
$E_{1}$ $\geq\frac{\big{(}a^{2}-ab\big{)}+\sqrt{(1-a^{2})}\big{(}\sqrt{1-a^{2}}-\sqrt{1-b^{2}}\big{)}}{2}$ (39)
$=\frac{1-ab-\sqrt{(1-a^{2})(1-b^{2})}}{2}$
$\geq\frac{1-ab-\sqrt{\big{(}1-\frac{a^{2}+b^{2}}{2}\big{)}^{2}}}{2}$
$=\frac{(a-b)^{2}}{4}\geq\frac{p}{4\mu r}\cdot r_{1}^{2}(a-b)^{2},$
where in the last step we use $r_{1}^{2}\leq\frac{\mu r}{p}$ as dictated in the $\mu$-incoherence condition.
On the other hand,
$E_{2}=E_{2}(a^{2})=1-\frac{a^{2}+\sqrt{(1-a^{2})^{2}+\lambda^{2}(1-a^{2})}}{\sqrt{1+\lambda^{2}}}.$
We have $E_{2}(0)=0$. Taking derivative with respect to $a^{2}$,
$E_{2}^{\prime}(a^{2})$ $=\frac{1}{\sqrt{1+\lambda^{2}}}\big{(}-1+\frac{2(1-a^{2})+\lambda^{2}}{2\sqrt{(1-a^{2})^{2}+\lambda^{2}(1-a^{2})}}\big{)}$
$=\frac{1}{\sqrt{1+\lambda^{2}}}\big{(}-1+\sqrt{\frac{4(1-a^{2})^{2}+4\lambda^{2}(1-a^{2})+\lambda^{4}}{4(1-a^{2})^{2}+4\lambda^{2}(1-a^{2})}}\big{)}\geq 0$
Therefore, $E_{2}\geq 0$ and $1-s^{\top}t\geq E_{1}+E_{2}\geq\frac{p}{8\mu r}\cdot r_{1}^{2}(a-b)^{2}$.
B.4 Proof of Lemma B.4
@@
We start the proof by exploring more geometric structure of $V^{\top}Re_{1}=\begin{bmatrix}t_{1}\\ t_{2}\\ \cdots\\ t_{r}\end{bmatrix}$. Denote $v_{1}-r_{1}=\lambda$, then $a^{2}(r_{1}-v_{1})^{2}=a^{2}\lambda^{2}$. Consider expansion of $Re_{1}$ on the incomplete basis of $Ve_{1},Ve_{2},\cdots,Ve_{r}$,
@@
$=t_{1}v_{1}+\sum_{i=2}^{r}t_{i}e_{1}^{\top}Ve_{i}+\alpha_{1}$
Thus,
@@
$\leq|\alpha_{1}|+\big{(}|1-t_{1}|+\sum_{i=2}^{r}|t_{i}|\big{)}\cdot\sqrt{\frac{\mu r}{p}},$ (40)
which will be crucial for later proof.
@@
$1=\|Re_{1}\|^{2}=\sum_{i=1}^{r}t_{i}^{2}+\|\alpha\|^{2}\geq\sum_{i=1}^{r}t_{i}^{2}+\alpha_{1}^{2}$ (41)
$\mu$-incoherence condition gives
@@
which will be extensively used later.
We consider several cases in the sequel:
@@
    $\geq\frac{\alpha_{1}^{2}}{2}\geq\frac{\lambda^{2}}{32}\geq\frac{a^{2}\lambda^{2}}{32}\frac{p}{\mu r}$
@@
  • Case 2: $\exists\ 2\leq k\leq r$ such that $|s_{k}|\geq\frac{1}{4r}\sqrt{\frac{p}{\mu r}}|\lambda|$. Recall that $\begin{bmatrix}s_{1}\\ s_{2}\\ \cdots\\ s_{r}\end{bmatrix}=U^{\top}Le_{1}$, i.e., $s_{k}=e_{1}^{\top}L^{\top}Ue_{k}$. Now consider $L^{\top}Ue_{k}=\begin{bmatrix}\widehat{s}_{1}\\ \widehat{s}_{2}\\ \cdots\\ \widehat{s}_{r}\end{bmatrix}$. Similarly as $\mathcal{E}_{r}$, we can define
@@
    Note that $\widehat{s}_{1}=e_{1}^{\top}L^{\top}Ue_{k}=s_{k}$, we have
    We have the decomposition of $\|L^{\top}Ue_{k}\|^{2}$ as follows,
    $\|L^{\top}Ue_{k}\|^{2}=\sum_{i=1}^{r}\widehat{s}_{i}^{2}=-\frac{b^{2}\widehat{s}_{1}^{2}}{1-b^{2}}+\frac{\widehat{s}_{1}^{2}}{1-b^{2}}+\sum_{i=2}^{r}\widehat{s}_{i}^{2}$ (43)
    Now because $L^{\top}Ue_{k}\in\widehat{\mathcal{E}}_{r}$, it holds that $\frac{\widehat{s}_{1}^{2}}{1-b^{2}}+\sum_{i=2}^{r}\widehat{s}_{i}^{2}\leq 1$.
    Plugging in (43) we have,
@@
    $\leq 1-\frac{a^{2}|\lambda|^{2}}{16r^{2}}\frac{p}{\mu r}$ (44)
    where the second inequality follows from $\widehat{s}_{1}=s_{k}$ and $a\leq b$, the last inequality follows from the assumption of Case 2 that $|s_{k}|\geq\frac{1}{4r}\sqrt{\frac{p}{\mu r}}|\lambda|$.
@@
    $1-\|L^{\top}Ue_{k}\|\geq 1-\sqrt{1-\frac{a^{2}|\lambda|^{2}}{16r^{2}}\frac{p}{\mu r}}\geq\frac{a^{2}|\lambda|^{2}}{32r^{2}}\frac{p}{\mu r}$
    Combining above we get $r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{a^{2}|\lambda|^{2}}{32r^{2}}\frac{p}{\mu r}$.
  • Case 3: $\exists\ 2\leq k\leq r$ such that $|s_{k}-t_{k}|\geq\frac{1}{4r}\sqrt{\frac{p}{\mu r}}|\lambda|$.
@@
    where the second inequality is by applying Lemma B.2 on $w_{s}$ and $w_{t}$.
@@
    $\geq\frac{|\lambda|^{2}}{32r^{2}}\frac{p}{\mu r}\geq\frac{a^{2}|\lambda|^{2}}{32r^{2}}\frac{p}{\mu r}$
    where the second inequality utilizes the assumption of Case 3 that $|s_{k}-t_{k}|\geq\frac{1}{4r}\sqrt{\frac{p}{\mu r}}|\lambda|$ and the last inequality uses $a^{2}\leq 1$.
  • Case 4: $t_{1}\leq 0$. Similar to the proof of Lemma B.3, we again exploit the geometric property of $s=U^{\top}Le_{1}$ and $t=V^{\top}Re_{1}$.
@@
    Recall the definition of $\mathcal{E}_{2}$ which is a ellipse centered at $(ab,0)$.
@@
    Observe that $\sqrt{(s_{1}+ab)^{2}+s_{2}^{2}}$ is the distance from a point on the ellipse to the center, thus should be no larger than the semi-major axis of the ellipse, which is $\sqrt{1-b^{2}}\leq 1-\frac{b^{2}}{2}$.
@@
    $s^{\top}t\leq 1-\frac{b^{2}}{2}\quad\forall s\in\mathcal{E}_{r},t\in\mathcal{T}_{r},t_{1}\leq 0$ (46)
@@
    $\forall\ 2\leq k\leq r,|s_{k}-t_{k}|<\frac{1}{4r}\sqrt{\frac{p}{\mu r}}|\lambda|,|s_{k}|<\frac{1}{4r}\sqrt{\frac{p}{\mu r}}|\lambda|.$
@@
    $\forall\ 2\leq k\leq r,|t_{k}|<\frac{1}{2r}\sqrt{\frac{p}{\mu r}}|\lambda|.$ (47)
    Also as the assumption of Case 1 doesn’t hold, we have $|\alpha_{1}|<\frac{1}{4}|\lambda|$. Reconsider the inequality (B.4), we can deduce that
    $|1-t_{1}|\geq\frac{1}{4}\sqrt{\frac{p}{\mu r}}|\lambda|$
@@
    $0\leq t_{1}\leq 1-\frac{1}{4}\sqrt{\frac{p}{\mu r}}|\lambda|$ (48)
@@
    $\|V^{\top}Re_{1}\|^{2}$ $=\sum_{i=1}^{r}t_{i}^{2}=t_{1}^{2}+\sum_{i=1}^{r-1}t_{i}^{2}$
    $\leq\big{(}1-\frac{1}{4}\sqrt{\frac{p}{\mu r}}|\lambda|\big{)}^{2}+(r-1)\cdot\big{(}\frac{1}{2r}\sqrt{\frac{p}{\mu r}}|\lambda|\big{)}^{2}$

### v2

@@ Appendix D Key Separation Lemma
D.1 Proof sketch
We first prove a simple discrepancy lemma D.2. Then, we do a more refined discrepancy analysis in Lemma D.3 and D.4 in the sense of norm discrepancy and direction discrepancy, respectively. Combining them, we get Lemma D.5, which is a special case of Lemma D.1 in which every singular values are the same. We then tackle the general case utilizing a technical lemma D.6, which is proved using linear programming technique.
Lemma D.2.
@@ D.1 Proof sketch
Lemma D.3.
Let $p>r\geq 2$, $\ U,V,L,R\in\mathcal{O}_{p,r}$, and assume they are all $\mu$-incoherent, denote $u=U^{\top}e_{1}$ and similarly for $v,l,r$. Also, denote $r_{1}$ as the first element of $r$. Suppose $u=ae_{1}$ and $l=be_{1}$, $0\leq a\leq b$ then we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{2\mu r}\cdot r_{1}^{2}(a-b)^{2}.$
Lemma D.4.
Under the same assumptions of Lemma D.3, we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{32\mu r^{2}}a^{2}(r_{1}-v_{1})^{2}.$
Combining the previous two lemmas D.3 and D.4 gives the next lemma which characterizes how the discrepancy between the first rows of $U,V,L,R$ impact the inner product $\langle U^{\top}L,V^{\top}R\rangle$.
Lemma D.5.
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq c\frac{p}{\mu r^{2}}\cdot(u^{\top}v-l^{\top}r)^{2}.$
Lemma D.5 solves the case where all singular values are identical for both $P$ and $Q$. Now we utilize the special case to prove the general case. We pause for a bit and state a technical lemma utilizing linear programming.
Lemma D.6.
Consider the optimization problem, which is essentially an LP problem.
@@
$\Omega=\Big{\{}X\in\mathbb{R}^{r\times r}\Big{|}\sum_{i=1}^{r}|X_{i,k}|\leq 1,\ \sum_{i=1}^{r}|X_{k,i}|\leq 1,\ \sum_{i=1}^{k}\sum_{j=1}^{k}X_{i,j}\leq k-\gamma_{k}\quad\forall 1\leq k\leq r\Big{\}}.$ (35)
@@
$\begin{bmatrix}0\\ x_{i}\end{bmatrix}=(I-e_{1}e_{1}^{\top})Ue_{i},\quad\begin{bmatrix}0\\ y_{i}\end{bmatrix}=(I-e_{1}e_{1}^{\top})Le_{i},$
@@
$U=\begin{bmatrix}a&0&\cdots&0\\ x_{1}&x_{2}&\cdots&x_{r}\end{bmatrix},\quad L=\begin{bmatrix}b&0&\cdots&0\\ y_{1}&y_{2}&\cdots&y_{r}\end{bmatrix}.$
Then we have
@@
$\begin{bmatrix}\frac{x_{1}}{\sqrt{1-a^{2}}},&x_{2},&x_{3},&\cdots,&x_{r}\end{bmatrix}\in\mathcal{O}_{p-1,r}.$
In other words, they form an incomplete basis of $\mathbb{R}^{p-1}$, hence
$(\frac{x_{1}^{\top}y_{1}}{\sqrt{1-a^{2}}})^{2}+\sum_{i=2}^{r}(x_{i}^{\top}y_{1})^{2}\leq\|y_{1}\|^{2}=1-b^{2}.$ (38)
@@
$\mathcal{E}_{r}=\Bigg{\{}\begin{bmatrix}s_{1}\\ s_{2}\\ \cdots\\ s_{r}\end{bmatrix}\in\mathbb{R}^{r}\Bigg{|}\frac{(s_{1}-ab)^{2}}{1-a^{2}}+\sum_{i=2}^{r}s_{i}^{2}\leq 1-b^{2}\Bigg{\}}.$
@@
Now we turn to $V^{\top}R$; it turns out it suffices to bound $e_{1}^{\top}V^{\top}Re_{1}$ to prove this lemma. In fact, using Lemma D.2 with $F=Ve_{1},G=Re_{1},r=1$, we get
$|e_{1}^{\top}V^{\top}Re_{1}|=\|F^{\top}G\|_{F}\leq\sqrt{1-(v_{1}-r_{1})^{2}}\leq 1-\frac{(v_{1}-r_{1})^{2}}{2}.$ (39)
@@
It turns out the structures of $\mathcal{E}_{r}$ and $\mathcal{T}_{r}$ are enough to prove the lemma. Specifically, we can prove for all $s\in\mathcal{E}_{r}$ and $t\in\mathcal{T}_{r}$, it holds that
$1-s^{\top}t\geq\frac{p}{8\mu r}\cdot r_{1}^{2}(a-b)^{2}.$ (40)
@@
$\begin{bmatrix}\widetilde{s}_{1}\\ \widetilde{s}_{2}\end{bmatrix}\in\mathcal{E}_{2},\begin{bmatrix}\tilde{t}_{1}\\ \tilde{t}_{2}\end{bmatrix}\in\mathcal{T}_{2},$ (42)
while by Cauchy-Schwarz Inequality, $s^{\top}t\leq\widetilde{s}_{1}\tilde{t}_{1}+\widetilde{s}_{2}\tilde{t}_{2}=\widetilde{s}^{\top}\tilde{t}$. Therefore we reduce to the case $r=2$.
@@
In fact, it suffices to rule out the case that the optimal $t^{\star}$ lies on the vertical line part of the border of $\mathcal{T}_{2}$. Suppose without loss of generality that $|t^{\star}_{1}|=1-\frac{(v_{1}-r_{1})^{2}}{2},\ t^{\star}_{2}\geq 0$, we immediate have the optimal $s^{\star}_{2}\geq 0$. Now define $\tilde{t}:=(t^{\star}_{1},\sqrt{1-t_{1}^{*2}})$. Let $\delta_{t}=\sqrt{1-t_{1}^{*2}}-t^{\star}_{2}>0$. We have $\tilde{t}\in\mathcal{T}_{2}$ and $s^{*T}\tilde{t}=s^{*T}t^{\star}+s^{\star}_{2}\delta_{t}>s^{*T}t^{\star}$, which is in contradiction with the optimality of $(s^{\star},t^{\star})$. Therefore, $t^{\star}$ does take the form of $\begin{bmatrix}\cos(\phi)\\ \sin(\phi)\end{bmatrix}$.
As $s^{\star}$ lies on the ellipse, it takes the form of $\begin{bmatrix}ab+\sqrt{(1-a^{2})(1-b^{2})}\cos{\theta}\\ \sqrt{1-b^{2}}\sin{\theta}\end{bmatrix}$. Therefore, we have
@@
$\leq\sqrt{\big{(}ab+\sqrt{(1-a^{2})(1-b^{2})}\cos{\theta}\big{)}^{2}+\big{(}\sqrt{1-b^{2}}\sin{\theta}\big{)}^{2}}$
$=\sqrt{a^{2}b^{2}+1-b^{2}+2ab\sqrt{1-a^{2}}\sqrt{1-b^{2}}\cos\theta-a^{2}(1-b^{2})\cos^{2}\theta}$
As $\frac{b\sqrt{1-a^{2}}}{a\sqrt{1-a^{2}}}=\sqrt{\frac{b^{2}-a^{2}b^{2}}{a^{2}-a^{2}b^{2}}}\geq 1$, the above display achieves maximum at $\cos\theta=1$ and hence
$s^{*T}t^{\star}\leq ab+\sqrt{1-a^{2}}\sqrt{1-b^{2}},$
which leads to
$1-s^{\top}t$ $\geq 1-s^{*T}t^{\star}$
$\geq 1-ab-\sqrt{1-a^{2}}\sqrt{1-b^{2}}$
$=\frac{(a-b)^{2}}{1-ab+\sqrt{1-a^{2}}\sqrt{1-b^{2}}}$
$\geq\frac{(a-b)^{2}}{2}$
$\geq\frac{p}{2\mu r}\cdot r_{1}^{2}(a-b)^{2}.$
where the last inequality is due to $|r_{1}|\leq\sqrt{\frac{\mu r}{p}}$ by the $\mu$-incoherence condition.
D.4 Proof of Lemma D.4
@@
We start the proof by exploring more geometric structure of $V^{\top}Re_{1}=\begin{bmatrix}t_{1}\\ t_{2}\\ \cdots\\ t_{r}\end{bmatrix}$. Denote $v_{1}-r_{1}=\lambda$, then $a^{2}(r_{1}-v_{1})^{2}=a^{2}\lambda^{2}$. Consider expansion of $Re_{1}$ on the incomplete basis $\{Ve_{1},Ve_{2},\cdots,Ve_{r}\}$,
@@
$=t_{1}v_{1}+\sum_{i=2}^{r}t_{i}e_{1}^{\top}Ve_{i}+\alpha_{1}.$
By the $\mu$-incoherence condition, we have
$\sqrt{\sum_{i=1}^{r}\big{(}e_{1}^{\top}Ve_{i}\big{)}^{2}}\leq\sqrt{\frac{\mu r}{p}}.$
Thus, by Cauchy-Schwartz inequality,
@@
$\leq|\alpha_{1}|+|1-t_{1}|\sqrt{\frac{\mu r}{p}}+\sqrt{\sum_{i=2}^{r}|t_{i}|^{2}}\cdot\sqrt{\frac{\mu r}{p}},$ (43)
, which will be crucial for the later proof. It is worth noting that the $\mu$-incoherence condition plays a nontrivial role in the proof exclusively through this equation.
@@
$1=\|Re_{1}\|^{2}=\sum_{i=1}^{r}t_{i}^{2}+\|\alpha\|^{2}\geq\sum_{i=1}^{r}t_{i}^{2}+\alpha_{1}^{2}.$ (44)
Also, $\mu$-incoherence condition yields
@@
, which will be extensively used later.
We consider five cases in the sequel:
@@
    $\geq\frac{\alpha_{1}^{2}}{2}\geq\frac{\lambda^{2}}{32}\geq\frac{a^{2}\lambda^{2}}{32}\frac{p}{\mu r},$
@@
  • Case 2: $\exists\ 2\leq k\leq r$ such that $|s_{k}|\geq\frac{1}{4}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|$. Recall that $\begin{bmatrix}s_{1}\\ s_{2}\\ \cdots\\ s_{r}\end{bmatrix}=U^{\top}Le_{1}$, i.e., $s_{k}=e_{1}^{\top}L^{\top}Ue_{k}$. Now consider $L^{\top}Ue_{k}=\begin{bmatrix}\widehat{s}_{1}\\ \widehat{s}_{2}\\ \cdots\\ \widehat{s}_{r}\end{bmatrix}$. Similarly as $\mathcal{E}_{r}$, we can define
@@
    We have the following decomposition of $\|L^{\top}Ue_{k}\|^{2}$,
    $\|L^{\top}Ue_{k}\|^{2}=\sum_{i=1}^{r}\widehat{s}_{i}^{2}=-\frac{b^{2}\widehat{s}_{1}^{2}}{1-b^{2}}+\frac{\widehat{s}_{1}^{2}}{1-b^{2}}+\sum_{i=2}^{r}\widehat{s}_{i}^{2}.$ (46)
    And because $L^{\top}Ue_{k}\in\widehat{\mathcal{E}}_{r}$, it holds that $\frac{\widehat{s}_{1}^{2}}{1-b^{2}}+\sum_{i=2}^{r}\widehat{s}_{i}^{2}\leq 1$.
    Plugging in (46) yields
@@
    $\leq 1-\frac{a^{2}|\lambda|^{2}}{16}\frac{p}{\mu r^{2}},$ (47)
    where the second inequality follows from $\widehat{s}_{1}=e_{1}^{\top}L^{\top}Ue_{k}=s_{k}$ and $0\leq a\leq b$, the last inequality follows from the assumption of Case 2 that $|s_{k}|\geq\frac{1}{4}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|$.
@@
    $1-\|L^{\top}Ue_{k}\|\geq 1-\sqrt{1-\frac{a^{2}|\lambda|^{2}}{16}\frac{p}{\mu r^{2}}}\geq\frac{a^{2}|\lambda|^{2}}{32}\frac{p}{\mu r^{2}}.$
    Combining above we get $r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{a^{2}|\lambda|^{2}}{32}\frac{p}{\mu r^{2}}$.
  • Case 3: $\exists\ 2\leq k\leq r$ such that $|s_{k}-t_{k}|\geq\frac{1}{4}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|$.
@@
    where the second inequality is by applying Lemma D.2 on $w_{s}$ and $w_{t}$.
@@
    $\geq\frac{|\lambda|^{2}}{32}\frac{p}{\mu r^{2}}\geq\frac{a^{2}|\lambda|^{2}}{32}\frac{p}{\mu r^{2}},$
    where the second inequality utilizes the assumption of Case 3 that $|s_{k}-t_{k}|\geq\frac{1}{4}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|$ and the last inequality uses $a^{2}\leq 1$.
  • Case 4: $t_{1}\leq 0$. Similar to the proof of Lemma D.3, we again exploit the geometric property of $s=U^{\top}Le_{1}$ and $t=V^{\top}Re_{1}$.
@@
    Recall the definition of $\mathcal{E}_{2}$ which is an ellipse centered at $(ab,0)$.
@@
    Observe that $\sqrt{(s_{1}-ab)^{2}+s_{2}^{2}}$ is the distance from a point on the ellipse to the center, thus should be no larger than the semi-major axis of the ellipse, which is $\sqrt{1-b^{2}}\leq 1-\frac{b^{2}}{2}$.
@@
    $s^{\top}t\leq 1-\frac{b^{2}}{2}\quad\forall s\in\mathcal{E}_{r},t\in\mathcal{T}_{r},t_{1}\leq 0.$ (49)
@@
    $\forall\ 2\leq k\leq r,\ |s_{k}-t_{k}|<\frac{1}{4}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|,\ |s_{k}|<\frac{1}{4}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|.$
@@
    $\forall\ 2\leq k\leq r,|t_{k}|<\frac{1}{2}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|.$ (50)
    Also, as the assumption of Case 1 doesn’t hold, we have $|\alpha_{1}|<\frac{1}{4}|\lambda|$. Revisiting the inequality (D.4), we can deduce that
    $|1-t_{1}|\geq\frac{1}{4}\sqrt{\frac{p}{\mu r}}|\lambda|.$
@@
    $0\leq t_{1}\leq 1-\frac{1}{4}\sqrt{\frac{p}{\mu r}}|\lambda|.$ (51)
@@
    $\|V^{\top}Re_{1}\|^{2}$ $=\sum_{i=1}^{r}t_{i}^{2}=t_{1}^{2}+\sum_{i=2}^{r}t_{i}^{2}$
    $\leq\big{(}1-\frac{1}{4}\sqrt{\frac{p}{\mu r}}|\lambda|\big{)}^{2}+(r-1)\cdot\big{(}\frac{1}{2}\sqrt{\frac{p}{\mu r^{2}}}|\lambda|\big{)}^{2}$